Darla is building a new desk. To make sure she had made a square corner, she measures 9ft. from the corner along one edge and 3 ft. fro mthe corner along the other edge. How long should the diagonal be between those two points if the corner is a right angle? I want to say 12 ft, but I think that is wrong, please help.
Answer by eperette(173)

Pythagorean Theorem: a^2 + b^2 = c^2
the diagonal is always longest side which is always c
(9)^2 + (3)^2 = c^2
81 + 9 = c^2
90 = c^2
sqr(90) = sqr (c^2)
sqr(9*10)= c
sqr(9) * sqr(10) = c
3sqr(10) = c
where sqr means square root